Evaluate the integral of (sec a tan a/(√e^(seca))) da from 0 to π/3


1
Expert's answer
2021-02-24T14:25:36-0500

"I" "=\int_{0}^{\pi/3} \dfrac{seca\times tana}{\sqrt{e^{seca}}}da"


"\rightarrow \int_{0}^{\pi/3} (e^{\dfrac{-seca}{2}})(seca\times tana )da"


Substitute, u= "e^{\dfrac{-seca}{2}}"


"Then, \dfrac{du}{da}=- \dfrac{seca\times tana}{2} \rightarrow da=- \dfrac{2}{seca\times tana}du"


On substituting, we get


"\int\dfrac{seca\times tana}{\sqrt{e^{seca}}}da=-2 \int e^udu" + C

"= -2 e^u +C"


Now put the value of u="e^{\dfrac{-seca}{2}}" in above equation,


"I= \mid-2e^{-\dfrac{seca}{2}}\mid_{0}^{\pi/3}"


On solving limit, we get

"I=\dfrac{2}{\sqrt e} - 2e^{-1}"


On simplification,

"I=e^{-1}(2\sqrt e-2)"


Hence the value of "\int_{0}^{\pi/3} \dfrac{seca\times tana}{\sqrt{e^{seca}}}da=" "e^{-1}(2\sqrt e-2)"


Approximation:

"I=0.4773024370823822"